OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of 5x^2 - 5x – 60 = 0.

OpenStudy (anonymous):

a=5, b=-5, c=-60 D=(-5)^2-4(5)(-60) =25+1200=1225>0 So Roots are real and unequal

OpenStudy (anonymous):

how many real roots are there?